OpenStudy (anonymous):

In a certain medium, sound travels at a speed of 7189 feet per second. How many meters per second does sound travel in this medium? Round to the nearest tenth.

OpenStudy (mathstudent55):

Convert feet into meters. The coversion factor is 1 ft = 0.3048 m. If you divide that equation by the left side you get\[\frac{ 1 ft }{ 1 ft } = \frac{ 0.3048 m }{ 1 ft }\] Then simplifiing the left side you get \[1 = \frac{ 0.3048 m }{ 1 ft }\] Since the fraction on the right equals 1, you may multiply any number by it without changing the number: \[7189 \frac{ ft }{ s } \times \frac{ 0.3048 m }{ 1 ft } \] The ft unitsa cancel out and you're left wit m/s